OpenStudy (anonymous):

Solve the system of equations: 3x-2y=-1 3x-4y=9

OpenStudy (anonymous):

you subtract i think its 2y- (-4y) which equals 6y and then subtract -1 -9 to get negative 10 so the answer is 6y=-10 and then divide so y=-5/3 Hopefully!! :p

OpenStudy (anonymous):

you can multiply the first equation by -2 and get \[-6x+4y=2\] \[3x-4y=9\] hten add to get \[-3x=11\] so \[x=-\frac{11}{3}\]

OpenStudy (anonymous):

then replace x by \[-\frac{11}{3}\] and solve for y you will get \[y=-5\]

OpenStudy (anonymous):

thank you